i need to some slabs for my garden which is 14ft 3in x 9ft 3in how many 450mm slabs would i need

thank you

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

14ft 3in x 9ft 3in
convert to mm
14ft 3in = (14*12)+3 =171 in
171*25.4 = 4343.4 mm
x 9ft 3in = (9*12)+3 =111 in
111*25.4= 2819.4 mm
Area of garden / area of each slab = number of tiles
Number of slabs = (2819.4*4343.4 )/(450*450)
=2820 slabs

Answer by ikleyn(53996)

.
i need to some slabs for my garden which is 14ft 3in x 9ft 3in how many 450mm slabs would i need
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


As it is worded in the post, the problem MAKES no SENSE.
In other words, it is NON-SENSICAL, like a compote of words.


The "solution" in the post by @mananth makes no sense, too.
In other words, it is non-sensical, too.


Many/several AI sites try to provide a "solution" to this "problem".
All these and such attempts are NON-SENSICAL.


Simply, for now, many AI solvers are not able to distinct meaningful problems from non-sensical gibberish.